# Multi-Agent Probabilistic Ensembles with Trajectory Sampling for Connected Autonomous Vehicles

### v1

@@ Abstract
Autonomous Vehicles (AVs) have attracted significant attention in recent years and Reinforcement Learning (RL) has shown remarkable performance in improving the autonomy of vehicles. In that regard, the widely adopted Model-Free RL (MFRL) promises to solve decision-making tasks in connected AVs (CAVs), contingent on the readiness of a significant amount of data samples for training. Nevertheless, it might be infeasible in practice and possibly lead to learning instability. In contrast, Model-Based RL (MBRL) manifests itself in sample-efficient learning, but the asymptotic performance of MBRL might lag behind the state-of-the-art MFRL algorithms. Furthermore, most studies for CAVs are limited to the decision-making of a single AV only, thus underscoring the performance due to the absence of communications. In this study, we try to address the decision-making problem of multiple CAVs with limited communications and propose a decentralized Multi-Agent Probabilistic Ensembles with Trajectory Sampling algorithm MA-PETS. In particular, in order to better capture the uncertainty of the unknown environment, MA-PETS leverages Probabilistic Ensemble (PE) neural networks to learn from communicated samples among neighboring CAVs. Afterwards, MA-PETS capably develops Trajectory Sampling (TS)-based model-predictive control for decision-making. On this basis, we derive the multi-agent group regret bound affected by the number of agents within the communication range and mathematically validate that incorporating effective information exchange among agents into the multi-agent learning scheme contributes to reducing the group regret bound in the worst case. Finally, we empirically demonstrate the superiority of MA-PETS in terms of the sample efficiency comparable to MFBL.
@@ I Introduction
Recently, there emerges significant research interest towards Connected Autonomous Vehicles (CAVs) with a particular emphasis on developing suitable Reinforcement Learning (RL)-driven controlling algorithms [2, 3] for the optimization of intelligent traffic flows [4], decision making [5], and control of AVs [6]. Notably, the research on Multi-Agent Reinforcement Learning (MARL) makes significant progress in capably learning complex tasks for CAVs, through effective interaction between agents and the environment. Existing MARL methods for CAVs can be classified as Model-Free Reinforcement Learning (MFRL) and Model-Based Reinforcement Learning (MBRL), the key differences of which lie in whether agents estimate an explicit environment model for the policy learning [7].
@@
Conventionally, MFRL, which relies on collected rewards on recorded state-action transition pairs, has been widely applied to model complex mixed urban traffic systems in multi-vehicle scenarios, showing excellent performance in various situations of autonomous driving [8, 9]. Typical examples of MFRL include MADDPG [10], COMA [11], QMIX [12], SVMIX[13]. However, the computational complexity of most MFRL algorithms grows exponentially with the number of agents. In order to solve training data scarcity-induced out-of-distribution (OOD) problems, MFRL is typically required to repetitively interact with the real-world to collect a sufficient amount of training data, which might be infeasible in practice and possibly lead to learning instability and huge overhead. On the contrary, due to the impressive sample efficiency, MBRL promises to solve CAVs issues[14] more capably and starts to attain some research interest [15, 16].
@@
Typically, MBRL is contingent on learning an accurate probabilistic dynamics model that can clearly distinguish between aleatoric and epistemic uncertainty [17], where the former is inherent to the system noise, while the latter stems from sample scarcity and contributes to solving the OOD problem to a certain extent. Afterwards, based on the learned dynamics model from the collected data, MBRL undergoes a planning and control phase by simulating model-consistent transitions and optimizing the policy accordingly. However, the asymptotic performance of MBRL algorithms has lagged behind state-of-the-art MFRL methods in common benchmark tasks, especially as the environmental complexity increases. In other words, although MBRL algorithms tend to learn faster, they often converge to poorer results [18, 19]. Fortunately, extending single-agent RL [20] to a multi-agent case through efficient communication can compensate for the learning deficiency to some extent. In that regard, despite the simplicity of assuming the existence of a central controller, it might be practically infeasible or cost-effective to install such a controller in many real-world scenarios. Meanwhile, it is often challenging to establish fully connected communication between all agents, where the required communication overhead can scale exponentially[8, 10]. Instead, a more complex communication-limited decentralized multi-agent MBRL for CAVs becomes feasible, and specifying a suitable protocol for cooperation between agents turns crucial.
@@
On the other hand, in order to theoretically characterise the sampling efficiency of RL, the concept of the regret bound emerges, which targets to theoretically measure the $T$-time-step difference between an agent’s accumulated rewards and the total reward that an optimal policy (for that agent) would have achieved. Without loss of generality, for an $H$-episodic RL environment with $S$ states and $A$ actions, contingent on Hoeffding inequality and Bernstein inequality, tabular upper-confidence bound (UCB) algorithms can lead to a regret bound of $\tilde{O}(\sqrt{H^{4}SAT})$ and $\tilde{O}(\sqrt{H^{3}SAT})$ respectively [21, 22], where $\tilde{\mathcal{O}}(\cdot)$ hides the logarithmic factors. On the other hand, for any communicating Markov decision process (MDP) with the diameter $D$, Jaksch et al. propose a classical confidence upper bound algorithm UCRL2 algorithm (abbreviation of Upper Confidence bound for RL) to achieve the regret bound $\tilde{O}(DS\sqrt{AT})$. In the literature, there only sheds a little light on the regret bound of collaborative MARL. In this paper, inspired by the regret bound of UCRL2, we investigate the regret bound of decentralized communication-limited MBRL, and demonstrate how communication among the multi-agents can be used to reduce the regret bound and boost the learning performance.
@@
In this paper, targeted at addressing the sample efficiency issue in a communication-limited multi-agent scenario, we propose a fully decentralized Multi-Agent Probabilistic Ensembles with Trajectory Sampling (MA-PETS) algorithm. In particular, MA-PETS could effectively exchange collected samples from individual agents to neighboring vehicles within a certain range, and extend the widely adopted single-agent Probabilistic Ensemble (PE) technique to competently reduce both aleatoric and epistemic uncertainty while fitting the multi-agent environmental transition. Furthermore, MA-PETS leverages computation-efficient Model Predictive Control (MPC)[23], which avoids gradient computations and exhibits appealing robustness to uncertainty, to produce suitable control actions from learned-model-based Trajectory Sampling (TS). Compared to the existing literature, the contribution of the paper can be summarized as follows:
@@
  • We formulate the decentralized multi-agent decision-making issue for CAVs as parallel time-homogeneous Markov Decision Process (MDP), and devise a sample efficient MBRL solution MA-PETS.
  • Within MA-PETS, we develop a multi-agent PE to learn the unknown environmental transition dynamics model from exchanged data samples and calibrate TS-based MPC for model-based decision-making.
@@
  • We analyze the group regret bound of MA-PETS on top of UCRL2, which theoretically demonstrates that multiple agents jointly exploring the state-action space in similar environments could communicate to discover the optimal policy faster than individual agents operating in parallel. On top of the concept of clique cover [24] of a communication undirected graph, our work mathematically verifies that in the worst case, the addition of more information from communication to a multi-agent algorithm still yields a sub-linear group regret concerning the number of agents and accelerates convergence. Thus, our theoretical derivation is completely different from [25], which takes the simple superposition of single-agent regret bounds.
@@ II-A Multi-Agent Reinforcement Learning of CAVs
Decision-making of CAVs in high-density mixed-traffic intersections is a challenging task. A set of different MARL approaches have been proposed to solve the decision-making of AVs at unsignalised intersections [27], so as to improve traffic efficiency and safety, and show remarkable performance in various CAV cases. For instance, within the actor-critic MFRL framework, [11] introduces the multi-agent actor-critic method COMA that employs a centralized critic for estimating the $\mathcal{Q}$-function, alongside decentralized actors to optimize policies. In a similar vein, [10] proposes using a central critic for centralized training coupled with multiple actors for distributed execution. However, a notable challenge arises as the number of agents increases or the action space expands, the computational complexity of MFRL algorithms like COMA escalates exponentially. To mitigate this computational burden, [12] proposes QMIX to involve the value decomposition of the joint value function into separate individual value functions for each agent. In addition, some researchers focus on the communication between agents. [13] proposes to adopt a stochastic graph neural network to capture the dynamic topological features of time-varying graphs while decomposing the value function. Nevertheless, the MFRL algorithm still suffers from high overhead due to the significant number of required calculations and samples.
@@ II-B Model-Based Reinforcement Learning
To address the sampling efficiency and communication overhead issues in MFRL, MBRL naturally emerges as an alternative solution. Unfortunately, MBRL suffers from performance deficiency, as it might fail to accurately estimate the uncertainty in the environment and characterize the dynamics model, which belongs to a critical research component in MBRL. For example, [28] proposes a DNN-based method that is, to some extent, qualified to separate aleatoric and epistemic uncertainty while maintaining appropriate generalization capabilities, while PILCO [29] marginalizes aleatoric and epistemic uncertainty of a learned dynamics model to optimize the expected performance. Another type of MBRL falls within the scope of Dyna-style methods [30], where additional data to improve the efficiency of the RL is generated from interactions between the RL and the virtual environment. [31] proposes an RL training algorithm MA-PPO that incorporates a prior model into PPO algorithm to speed up the learning process while maintaining the sampling efficiency. [19] takes advantage of Model Predictive Control (MPC) to optimize the RL agent’s behavior policy by predicting and planning within the modeled virtual environment at each training step. Until recently, [15] proposes an uncertainty-aware MBRL and verifies that it has competitive performance as the state-of-the-art MFRL. However, there is still little light shed on the multi-agent MBRL scenario, especially in the communication-limited case.
@@ II-C Regret bounds of Reinforcement Learning
Understanding the regret bound of online single-agent RL-based approaches to a time-homogeneous MDP has received considerable research interest. For example, [32] discusses the performance guarantees of a learned policy with polynomial scaling in the size of the state and action spaces. [33] introduces a UCRL algorithm and shows that its expected online regret in unichain MDPs is $O(\log T)$ after $T$ steps. Furthermore, [34] proposes the a UCRL2 algorithm which is capable of identifying an optimal policy through Extended Value Iteration (EVI), by conjecturing a set of plausible MDPs formed within the confidence intervals dictated by the Hoeffding inequality. Moreover, [34] demonstrates that the total regret for an optimal policy can be effectively bounded by $\tilde{O}(DS\sqrt{AT})$. Afterwards, many variants of UCRL2 have been proposed for the generation of tighter regret bounds. For instance, [35] proposes a UCCRL algorithm that derives sub-linear regret bounds $\tilde{O}\left(T^{\frac{2+\alpha}{2+2\alpha}}\right)$ with a parameter $\alpha$ for un-discounted RL in continuous state space. By using more efficient posterior sampling for episodic RL, [36] achieves the expected regret bound $\tilde{O}(\iota S\sqrt{AT})$ with an episode length $\iota$. [37] introduces a non-parametric tailored multiplier bootstrap method, which significantly reduces regret in a variety of linear multi-armed bandit challenges. Similar to the single-agent setting, agents in MARL attempt to maximize their cumulative reward by estimating value functions, and the regret bound can be analysed as well. For example, [38] proposes that multi-agent $Q$-learning with UCB-Hoeffding exploration through communication yields a regret $\mathcal{O}\left(\sqrt{IH^{4}SAT\iota}\right)$, where $I$ is the number of RL agents and $\iota:=$ $\log(SATI/p)$. Nevertheless, despite the progress on single-agent regret bound and the group regret bound for fully connected multi-agent cases [25], the group regret of MBRL in communication-limited multi-agent scenarios remains an unexplored area of research.
@@ III-A2 Model-Based Reinforcement Learning
An MBRL framework typically involves two phases (i.e., dynamics model learning and planning & control). In the dynamics model learning phase, each agent $i\in[I]$ estimates the dynamics model $\tilde{\mathcal{P}}^{(i)}$ from collected environmental transition samples by a continuous model (e.g., DNNs). Afterwards, based on the approximated dynamics model $\tilde{\mathcal{P}}^{(i)}$, the agent simulates the environment and makes predictions for subsequent action selection. In that regard, we can evaluate $w$-length action sequences $\mathbf{A}_{t:t+w-1}^{(i)}=\left(a_{t}^{(i)},\ldots,a_{t+w-1}^{(i)}\right)$ by computing the expected reward over possible state-action trajectories $\tau_{t:t+w-1}^{(i)}=\left({s}_{t}^{(i)},{a}_{t}^{(i)},\cdots,{s}_{t+w-1}^{(i)},{a}_{t+w-1}^{(i)},{s}_{t+w}^{(i)}\right)$, and optimize the policy accordingly. Mathematically, this planning & control phase can be formulated as

### v2

@@ Abstract
Autonomous Vehicles (AVs) have attracted significant attention in recent years and Reinforcement Learning (RL) has shown remarkable performance in improving the autonomy of vehicles. In that regard, the widely adopted Model-Free RL (MFRL) promises to solve decision-making tasks in connected AVs (CAVs), contingent on the readiness of a significant amount of data samples for training. Nevertheless, it might be infeasible in practice, possibly leading to learning instability. In contrast, Model-Based RL (MBRL) manifests itself in sample-efficient learning, but the asymptotic performance of MBRL might lag behind the state-of-the-art MFRL algorithms. Furthermore, most studies for CAVs are limited to the decision-making of a single AV only, thus underscoring the performance due to the absence of communications. In this study, we try to address the decision-making problem of multiple CAVs with limited communications and propose a decentralized Multi-Agent Probabilistic Ensembles (PEs) with Trajectory Sampling (TS) algorithm namely MA-PETS. In particular, to better capture the uncertainty of the unknown environment, MA-PETS leverages PE neural networks to learn from communicated samples among neighboring CAVs. Afterward, MA-PETS capably develops TS-based model-predictive control for decision-making. On this basis, we derive the multi-agent group regret bound affected by the number of agents within the communication range and mathematically validate that incorporating effective information exchange among agents into the multi-agent learning scheme contributes to reducing the group regret bound in the worst case. Finally, we empirically demonstrate the superiority of MA-PETS in terms of the sample efficiency comparable to MFBL.
@@ I Introduction
Recently, there has emerged significant research interest towards Connected Autonomous Vehicles (CAVs) with a particular emphasis on developing suitable Reinforcement Learning (RL)-driven controlling algorithms [2, 3] for the optimization of intelligent traffic flows [4], decision-making [5], and control of AVs [6]. Notably, these RL methods capably learn complex tasks for CAVs through effective interaction between agents and the environment, and existing RL works for CAVs can be classified as Model-Free Reinforcement Learning (MFRL) and Model-Based Reinforcement Learning (MBRL), the key differences of which lie in whether agents estimate an explicit environment model for the policy learning [7].
@@
Conventionally, MFRL, which relies on collected rewards on recorded state-action transition pairs, has been widely applied to model complex mixed urban traffic systems in multi-vehicle scenarios, showing excellent performance in various situations of autonomous driving [8, 9]. Typical examples of MFRL include MADDPG [10], COMA [11], QMIX [12], SVMIX[13]. However, the computational complexity of most MFRL algorithms grows exponentially with the number of agents. To solve training data scarcity-induced out-of-distribution (OOD) problems, MFRL is typically required to repetitively interact with the real world to collect a sufficient amount of training data, which might be infeasible in practice and possibly lead to learning instability and huge overhead. On the contrary, due to the impressive sample efficiency, MBRL promises to solve CAVs issues[14] more capably and starts to attain some research interest [15, 16].
@@
Typically, MBRL is contingent on learning an accurate probabilistic dynamics model that can clearly distinguish between aleatoric and epistemic uncertainty [17], where the former is inherent to the system noise, while the latter stems from sample scarcity and contributes to solving the OOD problem to a certain extent. Afterward, based on the learned dynamics model from the collected data, MBRL undergoes a planning and control phase by simulating model-consistent transitions and optimizing the policy accordingly. However, the asymptotic performance of MBRL algorithms has lagged behind state-of-the-art MFRL methods in common benchmark tasks, especially as the environmental complexity increases. In other words, although MBRL algorithms tend to learn faster, they often converge to poorer results [18, 19]. Consequently, the deployment of model-based strategies to intricate tasks often encounters substantial compounded errors, which may impede trajectory learning. However, by extending single-agent RL [20] to multi-agent contexts through efficient communication protocols, one can compensate for the learning deficiency to some extent[21, 22]. In that regard, despite the simplicity of assuming the existence of a central controller, it might be practically infeasible or cost-effective to install such a controller in many real-world scenarios. Meanwhile, it is often challenging to establish fully connected communication between all agents, where the required communication overhead can scale exponentially[8, 10]. Instead, a more complex communication-limited decentralized multi-agent MBRL for CAVs becomes feasible, and specifying a suitable protocol for cooperation between agents turns crucial.
@@
On the other hand, to theoretically characterize the sampling efficiency of RL, the concept of the regret bound emerges, which targets to theoretically measure the $T$-time-step difference between an agent’s accumulated rewards and the total reward that an optimal policy (for that agent) would have achieved. Without loss of generality, for an $H$-episodic RL environment with $S$ states and $A$ actions, contingent on Hoeffding inequality and Bernstein inequality, tabular upper-confidence bound (UCB) algorithms can lead to a regret bound of $\tilde{O}(\sqrt{H^{4}SAT})$ and $\tilde{O}(\sqrt{H^{3}SAT})$ respectively [23, 24], where $\tilde{\mathcal{O}}(\cdot)$ hides the logarithmic factors. On the other hand, for any communicating Markov decision process (MDP) with the diameter $D$, Jaksch et al. propose a classical confidence upper bound algorithm UCRL2 algorithm (abbreviation of Upper Confidence bound for RL) to achieve the regret bound $\tilde{O}(DS\sqrt{AT})$. In the literature, there only sheds little light on the regret bound of collaborative Multi-Agent Reinforcement Learning (MARL). In this paper, inspired by the regret bound of UCRL2, we investigate the regret bound of decentralized communication-limited MBRL, and demonstrate how communication among the multi-agents can be used to reduce the regret bound and boost the learning performance.
@@
In this paper, targeted at addressing the sample efficiency issue in a communication-limited multi-agent scenario, we propose a fully decentralized Multi-Agent Probabilistic Ensembles with Trajectory Sampling (MA-PETS) algorithm. In particular, MA-PETS could effectively exchange collected samples from individual agents to neighboring vehicles within a certain range, and extend the widely adopted single-agent Probabilistic Ensemble (PE) technique to competently reduce both aleatoric and epistemic uncertainty while fitting the multi-agent environmental transition. Furthermore, MA-PETS employs Model Predictive Control (MPC) [25, 26], due to its excellence in scenarios that necessitate the joint prediction and optimization of future behavior, to generate appropriate control actions from a learned-model-based Trajectory Sampling (TS) approach. Compared to the existing literature, the contribution of the paper can be summarized as follows:
@@
  • We formulate the decentralized multi-agent decision-making issue for CAVs as a parallel time-homogeneous Markov Decision Process (MDP). On this basis, we devise a sample-efficient MBRL solution MA-PETS which utilizes a multi-agent PE to learn the unknown environmental transition dynamics model from data exchanged within a limited range, and calibrate TS-based MPC for model-based decision-making.
@@
  • We analyze the group regret bound of MA-PETS, which is based on UCRL2 and fundamentally different from the approach in [22] that utilizes a mere superposition of single-agent regret bounds. By incorporating the concept of clique cover [27] for a limited-communication, undirected graph, we theoretically demonstrate that, even in the worst case, augmenting a multi-agent algorithm with additional communication information still results in a sub-linear group regret relative to the number of agents, and accelerates convergence. This validates that multiple agents jointly exploring the state-action space in similar environments could communicate to discover the optimal policy faster than individual agents operating in parallel.
@@ II-A Multi-Agent Reinforcement Learning of CAVs
Decision-making and planning are crucial components of CAV systems, with significant implications for enhancing the safety, driving experience, and efficiency of CAVs[3, 29]. The decision-making of CAVs in high-density mixed-traffic intersections belongs to one of the most challenging tasks and attracts significant research interest [30, 31, 32], towards improving traffic efficiency and safety. In this regard, Centralized Training and Decentralized Execution (CTDE) has become one of the most popular paradigms. For instance, within the actor-critic MFRL framework, [11, 10] proposes using a central critic for centralized training coupled with multiple actors for distributed execution. However, as the number of agents increases or the action space expands, the computational complexity of MFRL algorithms like MA-DDPG can rise exponentially, presenting a significant challenge. To mitigate this computational burden, [12] proposes QMIX to involve the value decomposition of the joint value function into separate individual value functions for each agent. In addition, [13] proposes to adopt a stochastic graph neural network to capture dynamic topological features of time-varying graphs while decomposing the value function. To avoid the large number of communication overheads that frequent information exchange incurs, [33] develops a consensus-based optimization scheme on top of the periodic averaging method, which introduces the consensus algorithm into Federated Learning (FL) for the exchange of a model’s local gradients.
Although these studies showcase the robust post-training performance of RL in real-time applications, the MFRL algorithm is still plagued by high overhead, stemming from the extensive computational and sampling requirements. This issue is particularly pronounced in scenarios such as CAVs[34], where data acquisition is challenging, interactions between agents and the environment are inefficient, and each iteration yields only a limited amount of data. Therefore, as data samples become scarce, the performance of MFRL can quickly degrade and become unstable.
@@ II-B Model-Based Reinforcement Learning
To address the sampling efficiency and communication overhead issues in MFRL, MBRL naturally emerges as an alternative solution. Unfortunately, MBRL suffers from performance deficiency, as it might fail to accurately estimate the uncertainty in the environment and characterize the dynamics model, which belongs to a critical research component in MBRL. For example, [35] proposes a DNN-based method that is, to some extent, qualified to separate aleatoric and epistemic uncertainty while maintaining appropriate generalization capabilities, while PILCO [36] marginalizes aleatoric and epistemic uncertainty of a learned dynamics model to optimize the expected performance. [19] takes advantage of MPC to optimize the RL agent’s behavior policy by predicting and planning within the modeled virtual environment at each training step. Another type of MBRL falls within the scope of Dyna-style methods [37], where additional data is generated from interactions between the RL and the virtual environment, thus improving the efficiency of the RL. [15] proposes an uncertainty-aware MBRL and verifies that it has competitive performance as the state-of-the-art MFRL. [8] proposes an RL training algorithm MAPPO that incorporates a prior model into PPO algorithm to speed up the learning process using current centralized coordination methods. However, centralized coordination methods face issues such as high resource-intensiveness, inflexibility, and annoying delays, since the central node must process a large amount of information and handle decision-making for the entire system. In addition, there is still little light shed on the multi-agent MBRL scenario, especially in the communication-limited case [38]. Considering the substantial communication overhead associated with centralized coordination methods, we mainly focus on the fully decentralized multi-agent MBRL algorithm under communication constraints.
@@ II-C Regret bounds of Reinforcement Learning
Understanding the regret bound of online single-agent RL-based approaches to a time-homogeneous MDP has received considerable research interest. For example, [39] discusses the performance guarantees of a learned policy with polynomial scaling in the size of the state and action spaces. [40] introduces a UCRL algorithm and shows that its expected online regret in unichain MDPs is $O(\log T)$ after $T$ steps. Furthermore, [41] proposes the a UCRL2 algorithm which is capable of identifying an optimal policy through Extended Value Iteration (EVI), by conjecturing a set of plausible MDPs formed within the confidence intervals dictated by the Hoeffding inequality. Moreover, [41] demonstrates that the total regret for an optimal policy can be effectively bounded by $\tilde{O}(DS\sqrt{AT})$. Afterward, many variants of UCRL2 have been proposed for the generation of tighter regret bounds. For instance, [42] proposes a UCCRL algorithm that derives sub-linear regret bounds $\tilde{O}\left(T^{\frac{2+\alpha}{2+2\alpha}}\right)$ with a parameter $\alpha$ for un-discounted RL in continuous state space. By using more efficient posterior sampling for episodic RL, [43] achieves the expected regret bound $\tilde{O}(\iota S\sqrt{AT})$ with an episode length $\iota$. [44] introduces a non-parametric tailored multiplier bootstrap method, which significantly reduces regret in a variety of linear multi-armed bandit challenges. Similar to the single-agent setting, agents in MARL attempt to maximize their cumulative reward by estimating value functions, and the regret bound can be analyzed as well. For example, [45] proposes that a specific class of online, episodic, tabular multi-agent Q-learning problems with UCB-Hoeffding exploration through communication yields a regret $\mathcal{O}\left(\sqrt{IH^{4}SAT\iota}\right)$, where $I$ is the number of RL agents and $\iota:=$ $\log(SATI/p)$. Nevertheless, despite the progress on single-agent regret bound and the group regret bound for fully connected multi-agent cases [22], the group regret of MBRL in communication-limited multi-agent scenarios remains an unexplored area of research. Consequently, we integrate dynamic graph theory, akin to that is found in [45], with an analysis of small blocks to explore the group regret bounds within a distinct MARL algorithm.
@@ III-A2 Model-Based Reinforcement Learning
An MBRL framework typically involves two phases (i.e., dynamics model learning, and planning & control). In the dynamics model learning phase, each agent $i\in[I]$ estimates the dynamics model $\tilde{\mathcal{P}}^{(i)}$ from collected environmental transition samples by a continuous model (e.g., DNNs). Afterward, based on the approximated dynamics model $\tilde{\mathcal{P}}^{(i)}$, the agent simulates the environment and makes predictions for subsequent action selection. In that regard, we can evaluate $w$-length action sequences $\mathbf{A}_{t:t+w-1}^{(i)}=\left(a_{t}^{(i)},\ldots,a_{t+w-1}^{(i)}\right)$ by computing the expected reward over possible state-action trajectories $\tau_{t:t+w-1}^{(i)}=\left({s}_{t}^{(i)},{a}_{t}^{(i)},\cdots,{s}_{t+w-1}^{(i)},{a}_{t+w-1}^{(i)},{s}_{t+w}^{(i)}\right)$, and optimize the policy accordingly. Mathematically, this planning & control phase can be formulated as